# Min-K%++: Improved Baseline for Detecting Pre-Training Data from Large Language Models

### v2

@@ 4.1 Setup
Benchmarks. We focus on two benchmarks (and the only two to our knowledge) for pre-training data detection, WikiMIA [28] and MIMIR [10]. WikiMIA is constructed by the authors of Min-Kand is the first benchmark for pre-training data detection. It consists of texts from Wikipedia events, and the training v.s. non-training data is determined by the timestamp (i.e., all events that happen after the release date of a LLM are guaranteed to be non-training data for that model). WikiMIA specifically groups data into splits according to the sentence length, intending to provide a fine-grained evaluation and isolate the effect of input length. It also considers two settings: original and paraphrased. The former assesses the detection of verbatim training texts, while the latter paraphrases the training texts (using ChatGPT) and evaluates on paraphrased inputs.
@@
Baselines. We consider 6 representative and state-of-the-art methods as our baselines, which are also featured in the benchmarking work of Duan et al. [10]. Except Min-K%, all methods consider sentence-level probability. Loss method [37] is a general technique that directly takes the loss as the score for detection. Reference method [4] (Ref) uses a reference LLM to calibrate the likelihood of the input. Zlib and Lowercase method [4] are also reference-based, except that they use zlib compression entropy and the likelihood of lowercased text as the reference, respectively. Neighbor method [21] perturbs the input sentence with masked language models to create “neighbors” and calibrate the loss of the input sentence with the average loss of the neighbor sentences. Lastly, as discussed earlier, Min-K[28] examines the token probability and aggregate the score through min-$k$selection; it is currently the best-performing method on WikiMIA.
Determining the $k$%. Our Min-K%++, just like Min-K%, selects the $k$of tokens with the minimum score to compute the overall score for the sentence. As existing benchmarks do not provide validation data, we sweep over 10, 20, …, 100 for $k$ and report the best result for both Min-Kand Min-K%++ for fair comparison. Later we show that Min-K%++ is robust to this hyperparameter in ablation study.
@@ 4.2 Results
WikiMIA results. Table 1 shows major results in terms of AUROC; for results on more models and TPR numbers, please see Appendix B Tables 5, 6, 7 and 8. We remark that Min-K%++ achieves significant improvements over existing methods. In the original setting, Min-K%++ on average outperforms the runner-up Min-Kby {10.5%, 12.3%, 6.2%} with inputs of length {32, 64, 128}, respectively. In the paraphrased setting, Min-K%++ is also the best-performing approach on average in all cases.
@@
MIMIR results. Table 2 shows the AUROC results; see Appendix B Table 9 for TPR results. Most numbers are taken from those reported by Duan et al. [10]. MIMIR is extremely challenging in that the training and non-training texts are maximally similar to each other since they are drawn from the same dataset. Nonetheless, Min-K%++ still improves upon existing methods in most cases. Averaged over 7 subsets, Min-K%++’s relative AUROC w.r.t. the powerful Min-Kis {–0.2%, +0.5%, +1.1%, +1.8%, +2.8%} on Pythia model with {160M, 1.4B, 2.8B, 6.9B, 12B} parameters, respectively. Extrapolating this trend, we anticipate Min-K%++’s effectiveness to be even more obvious with larger models. The most powerful method on MIMIR is the Reference approach [4] (Ref), the results of which are obtained by exhaustively trying out 8 different LLMs as the reference model and picking the best one [10] (note that inferencing with an extra LLM is expensive and may not always be feasible). In contrast, our Min-K%++ does not rely on a reference model, yet provides competitive performance that is on par with Ref. Meanwhile, we remark that Min-K%++ achieves new SOTA results among all the 5 reference-free methods out of the 6 total baselines. Also, in Appendix B Table 9, Min-K%++ outperforms all baselines when it comes to TPR performance.
@@ 4.3 Ablation study
Ablation on the hyperparameter $k$%. $k$ determines what percent of tokens with minimum scores are chosen to compute the final score. From Table 3, it is obvious that Min-K%++ is robust to the choice of $k$, with the best and the worst result being 84.8% and 82.1% (a variation of 2.7%), respectively. Min-Kis a little more sensitive to $k$: the variation between the best (68.0%) and the worst result (63.6%) is 4.4%, slightly larger than that of Min-K%++.
Decomposing the contribution of calibration factors. Recall that our method can be interpreted as calibrating the log probability $\log p(x_{t}|x_{<t})$ with two calibrationg factors $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ (Equation 5). In Table 3, we decompose the effect of $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. Specifically, starting from the raw log probability $\log p(x_{t}|x_{<t})$ (which degenerates to the design of Min-K%), we gradually incorporate $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ into the score computation, until we reach the final formulation of our Min-K%++. For example, when $\log p(x_{t}|x_{<t})$ and $\sigma_{x_{<t}}$ are included (marked by ✓in Table 3), the token-wise score becomes $\frac{\log p(x_{t}|x_{<t})}{\sigma_{x_{<t}}}$. From the results, we see that calibrating the log probability with either $\mu_{x{<t}}$ or $\sigma_{x{<t}}$ alone already leads to 9.3% and 7.0% performance boosts over using raw log probability (i.e., Min-K%). Combining them together, which leads to the formulation specified by Min-K%++, takes advantage from both factors and results in a larger improvement of 16.8%. Such observation again validates the proposed method.
@@
Table 3: Decomposing the token-wise score of Min-K%++ in Equation 5. The first row corresponds to Min-K%, while the last row corresponds to the final design of our Min-K%++.
@@ 4.4 An online detection setting
Setup. Current benchmarks fail to simulate the online detection setting, since each whole input is either pure training text or pure non-training text. To address this, we adapt WikiMIA to construct an online version of the benchmark. Concretely, each input text is created by concatenating a training text at the end of a non-training text, closely simulating the representative scenario discussed above. Both the training and non-training text have random length, varying among {32, 64, 128}. In this online setting, the prediction on each part of the input, instead of on the whole input, is of interests. Therefore, we split each input into chunks with a length of 32. Methods will be operating on all chunks, and the performance is evaluated on the chunk level. Essentially, this setup simulates using a non-overlapping sliding window of size 32 to sweep over the texts and detecting whether each part within the sliding window is training data or not. See Appendix B Figure 3 for visual illustration.
@@ 5 Related Work
Membership Inference Attacks. MIA has long been a security and privacy-related topic. Given a target model and a (arbitrary) target input, MIA aims to infer whether the input has been used to train the target model [29, 37]. It has been extensively studied in both vision [5, 6, 8, 38] and language domains [4, 34, 21]. Besides investigation on advanced methodologies, MIA also has profound implications in quantifying privacy risks [23], measuring memorization [7], helping with privacy auditing [30, 36], and detecting test-set contamination [26] and copyrighted contents [22, 11].
Pre-training data detection for LLMs. Despite being an instance of MIA (the general definition and evaluation metrics remain the same), this problem poses unique challenges compared with conventional MIA settings [28, 10]. First, certain assumptions such as no access to subsets of known training data (as nowadays many LLMs’ pre-training corpus is nonpublic) make some early methods that need to train shadow models on the same distribution of the target model impractical. Second, characteristics of LLM pre-training (e.g., few training epochs, large-scale training) inherently makes MIA much more challenging. Lastly, most existing works on MIA against LLMs target the fine-tuning stage [24, 12, 21], which cannot transfer to the pre-training stage for the same reasons.
As a result, there has been growing interests in pre-training data detection recently, though it is still largely underexplored. Shi et al. [28], to our knowledge, is the first to investigate this problem. They contribute the WikiMIA benchmark and propose the Min-Kmethod as a simple yet powerful baseline. Duan et al. [10] construct the MIMIR benchmark, perform systematic evaluation of prior approaches, and analyze the challenges of pre-training data detection for LLMs. In this work we propose a novel method that outperforms Min-Kand achieve superior performances over existing reference-free methods on both benchmarks. There are also works that specifically investigate copy-righted content detection, e.g., identifying memorized books [11].
@@ 6 Conclusion and Discussion
In this work, we propose Min-K%++ as a novel method for pre-training data detection for LLMs. Motivated by our insight that maximum likelihood training implicitly optimizes the Hessian trace of likelihood over training data, we design our method to measure how sharply peaked the likelihood is around an input under the discrete distribution of the LLM. It consistently achieves superior performances on two existing benchmarks and in various settings, which establishes a solid baseline for future studies. We hope that our method, along with our theoretical and empirical analysis, can motivate more research upon the pre-training data detection problem.
@@
Future work. Even with the proposed Min-K%++, detection on MIMIR is still challenging given the limited distribution shifts between training and non-training data. Understanding why and how exactly distribution shifts affect detection remains an important direction for future study. Meanwhile, our theoretical insight discussed in Section 3.2 is around general maximum likelihood training and has strong potential to generalize to other modalities, e.g., generative image modeling [27]. Investigating whether the concept of our method can work for other applications would also be meaningful.
@@
Appendix A Further Discussion on Min-K%++
Here we discuss that the numerator of Min-K%++’s formulation in Equation 5 shares a similar form to an approximation of the negative Hessian trace. This further justifies our design since we have shown that the Hessian trace is implicitly optimized over training data during maximum likelihood training.
Proposition 1.
The negative Hessian trace $-\text{tr}(H_{f}(x))$ of a function $f$ can be approximated with $f(x)-\mathbb{E}_{n}f(x+n)$, with mild constraints placed on $n$.
1 is proved by Mitchell et al. [25], and we refer readers to their work for details. If we denote the log likelihood function to be $f$, i.e., $\log p(\cdot|x_{<t})=f(\cdot)$, then the numerator of Min-K%++ token score, $\log p(x_{t}|x_{<t})-\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$, can be written as $f(x_{t})-\mathbb{E}_{z}f(z)$, where $x_{t}$ is the $t$-th token of the input, and $z$ enumerates all tokens in the vocabulary. This follows a similar form to $f(x)-\mathbb{E}_{n}f(x+n)$, meaning that (in a loose sense) the numerator in Equation 5 is approximating the negative Hessian trace, a statistic that is optimized over training instances and thus can be informative for training data detection.
Appendix B Additional Results

### v3

@@ 4.1 Setup
Benchmarks. We focus on two benchmarks (and the only two to our knowledge) for pre-training data detection, WikiMIA [29] and MIMIR [10]. WikiMIA is the first benchmark for pre-training data detection, which consists of texts from Wikipedia events. The training v.s. non-training data is determined by the timestamp. WikiMIA specifically groups data into splits according to the sentence length, intending to provide a fine-grained evaluation. It also considers two settings: original and paraphrased. The former assesses the detection of verbatim training texts, while the latter paraphrases the training texts (using ChatGPT) and evaluates on paraphrased inputs.
@@
Baselines. We consider 6 representative and state-of-the-art methods as our baselines, which are also featured in the benchmarking work of Duan et al. [10]. Except Min-K%, all methods consider sentence-level probability. Loss method [39] is a general technique that directly takes the loss as the score for detection. Reference method [4] (Ref) uses a reference LLM to calibrate the likelihood of the input. Zlib and Lowercase method [4] are also reference-based, except that they use zlib compression entropy and the likelihood of lowercased text as the reference, respectively. Neighbor method [23] perturbs the input sentence with masked language models to create “neighbors” and calibrate the loss of the input sentence with the average loss of the neighbor sentences. Lastly, Min-K[29] examines the exact token probability and averages a subset of minimum token scores over the input; it is currently the best-performing method on WikiMIA.
Determining the $k$%. Our Min-K%++ has a hyperparameter $k$ (Equation 4). As existing benchmarks do not provide validation data, we sweep over 10, 20, …, 100 for $k$ and report the best result. For fair comparison, we also report the best possible performance of other methods. Later we show that Min-K%++ is robust to this hyperparameter.
@@ 4.2 Results
WikiMIA results. Table 1 shows major results in terms of AUROC; for results on more models and TPR numbers, please see Appendix C Tables 5, 6, 7 and 8. We remark that Min-K%++ achieves significant improvements over existing methods. In the original setting, Min-K%++ on average outperforms the runner-up Min-Kby {10.5%, 12.3%, 6.2%} with inputs of length {32, 64, 128}, respectively. In the paraphrased setting, Min-K%++ is also the best-performing approach on average in all cases.
@@
MIMIR results. Table 2 shows the AUROC results; see Appendix C Table 9 for TPR results. Most numbers are taken from those reported by Duan et al. [10]. MIMIR is extremely challenging in that the training and non-training texts are maximally similar to each other since they are drawn from the same dataset. Nonetheless, Min-K%++ still improves upon existing methods in most cases. Averaged over 7 subsets, Min-K%++’s relative AUROC w.r.t. the powerful Min-Kis {–0.2%, +0.5%, +1.1%, +1.8%, +2.8%} on Pythia model with {160M, 1.4B, 2.8B, 6.9B, 12B} parameters, respectively. Extrapolating this trend, we anticipate Min-K%++’s effectiveness to be even more obvious with larger models. The most powerful method on MIMIR is the Reference approach [4] (Ref), the results of which are obtained by exhaustively trying out 8 different LLMs as the reference model and picking the best one [10] (note that inferencing with an extra LLM is expensive and may not always be feasible). In contrast, our Min-K%++ does not rely on a reference model, yet provides competitive performance that is on par with Ref. Meanwhile, we remark that Min-K%++ achieves new SOTA results among all the 5 reference-free methods out of the 6 total baselines. Also, in Appendix C Table 9, Min-K%++ outperforms all baselines when it comes to TPR performance.
@@ 4.3 Ablation study
Ablation on the hyperparameter $k$%. $k$ determines what percent of token sequences with minimum scores are chosen to compute the final score. From Table 3, it is obvious that Min-K%++ is robust to the choice of $k$, with the best and the worst result being 84.8% and 82.1% (a variation of 2.7%), respectively. Min-Khas a similar hyperparameter but is more sensitive to it: the variation between the best (68.0%) and the worst result (63.6%) is 4.4%, slightly larger than that of Min-K%++.
Decomposing the contribution of calibration factors. Recall that our method can be interpreted as calibrating the log probability $\log p(x_{t}|x_{<t})$ with two calibrationg factors $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ (Equation 3). In Table 3, we decompose the effect of $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. Specifically, starting from the raw log probability $\log p(x_{t}|x_{<t})$, we gradually incorporate $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ into the score computation, until we reach the final formulation of our Min-K%++. For example, when $\log p(x_{t}|x_{<t})$ and $\sigma_{x_{<t}}$ are included (marked by ✓in Table 3), the token-wise score becomes $\frac{\log p(x_{t}|x_{<t})}{\sigma_{x_{<t}}}$. From the results, we see that calibrating the log probability with either $\mu_{x{<t}}$ or $\sigma_{x{<t}}$ alone already leads to 9.3% and 7.0% performance boosts over using raw log probability. Combining them together, which leads to the formulation specified by Min-K%++, takes advantage from both factors and results in a larger improvement of 16.8%. Such observation again validates the proposed method.
@@
Table 3: Decomposing the token-wise score of Min-K%++ in Equation 3. The last row corresponds to the final design of our Min-K%++.
@@ 4.4 An online detection setting
Setup. Current benchmarks fail to simulate the online detection setting, since each whole input is either pure training text or pure non-training text. To address this, we adapt WikiMIA to construct an online version of the benchmark. Concretely, each input text is created by concatenating a training text at the end of a non-training text, closely simulating the representative scenario discussed above. Both the training and non-training text have random length, varying among {32, 64, 128}. In this online setting, the prediction on each part of the input, instead of on the whole input, is of interests. Therefore, we split each input into chunks with a length of 32. Methods will be operating on all chunks, and the performance is evaluated on the chunk level. Essentially, this setup simulates using a non-overlapping sliding window of size 32 to sweep over the texts and detecting whether each part within the sliding window is training data or not. See Appendix C Figure 3 for visual illustration.
@@ 5 Related Work
Membership Inference Attacks. MIA has long been a security and privacy-related topic. Given a target model and a target input, MIA aims to infer whether the input has been used to train the target model [30, 39]. It has been extensively studied in both vision [5, 6, 8, 40] and language domains [4, 36, 23]. Besides investigations on advanced methodologies, MIA also has profound implications in quantifying privacy risks [25], measuring memorization [7], helping with privacy auditing [31, 38], and detecting test-set contamination [27] and copyrighted contents [24, 11].
Pre-training data detection for LLMs. Despite being an instance of MIA (the general definition and evaluation metrics remain the same), this problem poses unique challenges compared with conventional MIA settings [29, 10]. For example, many early MIA methods need to train shadow models on the same distribution as the target model, which is no longer practical as many LLMs’ pre-training corpus is nonpublic. Second, characteristics of LLM pre-training (e.g., few training epochs, large-scale training) inherently makes MIA much more challenging. Lastly, most existing works on MIA against LLMs target the fine-tuning stage [26, 12, 23], which cannot transfer to the pre-training stage for the same reasons.
As a result, there has been growing interests in pre-training data detection recently, though it is still largely underexplored. Shi et al. [29], to our knowledge, is the first to investigate this problem. They contribute the WikiMIA benchmark and propose the Min-Kmethod. Duan et al. [10] perform systematic evaluation with the constructed MIMIR benchmark and analyze the challenges. In this work we propose a novel method that outperforms Min-Kand achieve superior performances over existing reference-free methods on both benchmarks. There are also works that specifically investigate copy-righted content detection, e.g., identifying memorized books [11].
@@ 6 Conclusion and Discussion
In this work, we propose Min-K%++ as a novel method for pre-training data detection for LLMs. Motivated by our insight that training data tends to be local maximum or locates near local maximum along input dimensions, we design our method to examine whether the input forms the mode or has relatively high probability under the conditional categorical distribution of the LLM. It consistently achieves superior performances on two existing benchmarks and in various settings, which establishes a solid baseline for future studies. We hope that our method, along with our theoretical and empirical analysis, can motivate more research upon the pre-training data detection problem.
@@
Future work. Even with the proposed Min-K%++, detection on MIMIR is still challenging given the limited distribution shifts between training and non-training data. Understanding why and how exactly distribution shifts affect detection remains an important direction for future study. Meanwhile, as aforementioned, our theoretical insight discussed in Section 3.1 has potential to generalize to other modalities and models. Investigating whether the concept of our method can work for other applications would also be meaningful.
@@
Appendix A Existing Assets
The two existing assets used in this work are the WikiMIA benchmark [29] and the MIMIR benchmark [10]. Please see their information below.
Appendix B Broader Impacts
As our work proposes an advanced pre-training data detection method, we will discuss societal impacts around general pre-training data detection techniques.
Positive impacts. Pre-training data detection has positive implications for several serious issues. For example, it can help with privacy auditing [29] by detecting whether a data instance has indeed been forgot or unlearned by the model. It can also help content creators or media see whether their copyrighted contents have been illegally used to train LLMs [14, 20].
Negative impacts. Pre-training data detection, as an instance of membership inference attack, may be thought to exhibit negative impacts such as stealing intellectual property (proprietary training data) and exposing the sensitive information encoded in training samples [39]. However, such risks are much less profound and practical than data extraction attacks [8]. For example, the proprietary training corpus of big companies will most likely be kept in-house, which is nearly impossible for attackers to access in the first place, not to mention to perform the subsequent membership inference attack. Therefore, from a practical perspective, we posit that the risks of pre-training data detection as an “attack” would not pose severe negative societal impacts.
Appendix C Additional Results